[The sum of the ages of Sam and Thomas is 27 years. Seven years ago Thomas was 3 years more than one-fourth as old as Sam. Find their present ages.]

Answer by TimothyLamb(4379)

x = sam age
y = thomas age
---
x + y = 27
y - 7 = (1/4)(x - 7) + 3
---
put the system of linear equations into standard form
---
x + y = 27
y - 7 = (1/4)(x - 7) + 3
y - 7 = 0.25x - 7/4 + 3
-0.25x + y = 10 - 7/4
-0.25x + y = 40/4 - 7/4
-0.25x + y = 33/4
---
x + y = 27
-0.25x + y = 8.25
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = sam age = 15
y = thomas age = 12
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php